Project ABC has an initial cost of $1,000 and generates cash inflow of $2,000 at the end of year 1. Project QRS has an initial cost of $10,000 and has annual cash inflows of $4,400, $4,840, and $5,324 in year 1-3. Assuming a 10% discount rate, calculate the net present value and profitability index for both projects. Which project is better in terms of NPV? Which project is better in terms of PI?
Answer #1

Project ABC gnitiof Cost = 1000 Cash in flow at year 1= 2000 1000 2000 NPV 2000 1-10 - 1000 = $ 218.18 Profitability Index =
Answer #2

a) Using NPV, Project QRS is better since it has the higher NPV

b) Using PI, Project ABC is better since it has a higher PI


a) Calculate NPV

ABC: Using finl calculator, C0=-$1000, C1=$2000, I/Y = 10%, CMPT NPV=$818.18

QRS: Using finl calculator, C0=-$10,000, C1=$4,400, C2=$4,840, C3=$5,324, I/Y=10% CMPT NPV=$2000


b) Calculate PI

ABC: NPV/initial investment = $818.18/$1000 = 0.82

QRS: NPV/initial investment = $2000/$10,000 = 0.2


answered by: anonymous